OpenStudy (anonymous):

write a formula for the linearization L(x) of F(x) using this information: F(x) = -x2 /2 F '(-2) = 2

OpenStudy (anonymous):

i assume you want the linearization at x = -2

OpenStudy (anonymous):

and that your function is \[F(x)=-\frac{x^2}{2}\]

OpenStudy (anonymous):

in which case you need a point and a slope, and you use the point-slope formula: the point is \[(-2,F(-2))\] in other words \[(-2,-2)\] and \[m=F'(-2)=2\] so the linearization is \[y-(-2)=2(x-(-2))\] \[y+2=2x+4\] \[y=2x+2\] or even \[L(x)=2x+2\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw